OpenStudy (anonymous):

Solve for x: 1 over 2 plus 4 over 2 x equals quantity x plus 4 over 10 x = –4 x = 5 x = –4 and x = 5 No Solution i will reward medal

OpenStudy (anonymous):

\[ \frac{1}{2} + \frac{4x}{2} = x + \frac{4}{10}\] Like this?

OpenStudy (anonymous):

the last part is \[\frac{ x+4 }{ 10 }\]

OpenStudy (anonymous):

So whats the first step?

OpenStudy (anonymous):

oh and switch the 4x and 2

OpenStudy (anonymous):

so it is \[\frac{ 1 }{ 2 }+\frac{ 4 }{ 2x }=\frac{ x+4 }{ 10 }\]

OpenStudy (anonymous):

and the first step is to find a common denominator.and after that i get confused

OpenStudy (anonymous):

@Knivez

OpenStudy (anonymous):